OpenStudy (anonymous):

could someone please help me? Solve the system of equations. 5x - y = 3 and -10x + 2y = -6 (0, -3) (-3, 0) (3, 12) no solution infinitely many solutions

OpenStudy (campbell_st):

If you write the lines in slope intercept form it may help y = 5x - 3 2y = 10 - 6 (divide each term by 2 to get the slope intercept form) what do you notice...?

OpenStudy (campbell_st):

oops should be 2y = 10x - 6

OpenStudy (anonymous):

uum y=5-3? oooh y=2?

OpenStudy (anonymous):

im confused....

OpenStudy (anonymous):

addition/elimination 5x - y = 3---->(2)5x - y = 3 -10x + 2y = - 6 -------------- 10x - 2y = 3 (result of multiplying by 2) -10x + 2y = 6 -------------add 0x + 0y = 9 This system of equations have no solution because there are no values of x and y.

OpenStudy (anonymous):

ok thank you